OpenStudy (clongoria2):

Help please! Medal and Fan for answer! Bruce has a bottle that contains 60% of lemon juice and the rest water. The bottle has 1 liter of water. Part A: Write an equation in one variable that can be used to find the total number of liters of lemon juice and water in the bottle. Define the variable used in the equation. (5 points) Part B: How many liters of lemon juice are present in the bottle? Show your work. (5 points)

OpenStudy (gorv):

let total mixture = x

OpenStudy (gorv):

lemon= x*60/100=0.6x water = x-0.6x=0.4x

OpenStudy (gorv):

but water= 1 litre so 1 litre = 0.4x x=1/0.4 x=10/4=5/2=2.5 litre

OpenStudy (gorv):

so total mixture = 2.5 ltr water= 1 litre lemon= 2.5-1=1.5 litre

OpenStudy (clongoria2):

Thank you!!!! I actually understand how you did it. @gorv Medal for you!

OpenStudy (gorv):

welcm:)